# Pareto Control Barrier Function for Inner Safe Set Maximization Under Input Constraints

### v1

@@ III-C Pareto Control Barrier Function (PCBF)
Figure 3: The left subfigure visualizes the inner safe sets obtained by the LCCBF (blue solid line) and PCBF (orange solid line) methods, alongside the infinite-time viability kernel (black dashed line) computed using HJ reachability. The region enclosed by the black dashed line represents the true safe states. The right subfigure depicts the evolution of the inner safe sets (green solid lines) for both PCBF and LCCBF methods throughout the training process. During the initialization phase, both methods yield identical results as they both employ the linear combination loss initially.
Algorithm 1 Pareto Control Barrier Function (PCBF)
1:Initial parameter $\theta_{0}$, learning rate $\eta$, hyperparameters $\beta$, $\epsilon_{\text{ub}}$, $\epsilon_{\text{lb}}$, regularization weight $\gamma$, max iterations $T_{\text{max}}$, initialization iterations $T_{\text{init}}$, convergence threshold $\epsilon$, initial state $x_{e}$, covariance matrix $\Sigma$ used for Gaussian sampling.
2:Optimal parameter $\theta^{*}$
3:$t\leftarrow 0$, $\theta_{t}\leftarrow\theta_{0}$
4:// Gaussian sampling around $x_{e}$
5:$\mathbb{X}_{\mathcal{N}}\sim\mathcal{N}(x_{e},\Sigma)$
6:// Initialization Phase
7:for $t=0$ to $T_{\text{init}}$ do
8:     $\theta_{t+1}\leftarrow\theta_{t}-\eta(\lambda_{1}\nabla\mathcal{L}_{\text{feas}}(\theta_{t})+\lambda_{2}\nabla\mathcal{L}_{\text{vol}}(\theta_{t}))$
9:end for
10:// Parameter Update Phase
11:for $t=T_{\text{init}}$ to $T_{\text{max}}$ do
12:     if $\mathcal{L}_{\text{vol}}(\theta_{t})>\beta\mathcal{L}_{\text{feas}}(\theta_{t})+\epsilon_{\text{ub}}$ then
13:         $\lambda\leftarrow\max\left(\min\left(\frac{((1+\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\nabla\mathcal{L}_{\text{vol}}(\theta_{t}))^{T}}{\|(1+\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\nabla\mathcal{L}_{\text{vol}}(\theta_{t})\|_{2}^{2}}\right.\right.$
14:                  $\left.\left.\cdot\nabla\mathcal{L}_{\text{feas}}(\theta_{t}),1\right),0\right)$
15:         $d_{t}\leftarrow(\lambda+\lambda\beta-1)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\lambda\nabla\mathcal{L}_{\text{vol}}(\theta_{t})$
16:     else if $\beta\mathcal{L}_{\text{feas}}(\theta_{t})+\epsilon_{\text{lb}}\leq\mathcal{L}_{\text{vol}}(\theta_{t})\leq\beta\mathcal{L}_{\text{feas}}(\theta_{t})+\epsilon_{\text{ub}}$ then
17:         $\lambda\leftarrow\max\left(\min\left(\frac{(\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\nabla\mathcal{L}_{\text{vol}}(\theta_{t}))^{T}\nabla\mathcal{L}_{\text{feas}}(\theta_{t})}{\|\nabla\mathcal{L}_{\text{vol}}(\theta_{t})-\nabla\mathcal{L}_{\text{feas}}(\theta_{t})\|_{2}^{2}},\right.\right.$
[…]
19:         $d_{t}\leftarrow-\lambda\nabla\mathcal{L}_{\text{vol}}(\theta_{t})-(1-\lambda)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})$
20:     else
21:         $\lambda\leftarrow\max\left(\min\left(\frac{(\nabla\mathcal{L}_{\text{vol}}(\theta_{t})+(1-\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t}))^{T}}{\|\nabla\mathcal{L}_{\text{vol}}(\theta_{t})+(1-\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})\|_{2}^{2}}\right.\right.$
22:                  $\left.\left.\cdot\nabla\mathcal{L}_{\text{feas}}(\theta_{t}),1\right),0\right)$
23:         $d_{t}\leftarrow(\lambda-\lambda\beta-1)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})+\lambda\nabla\mathcal{L}_{\text{vol}}(\theta_{t})$
24:     end if
25:     $\theta_{t+1}\leftarrow\theta_{t}+\eta(d_{t}-\gamma\nabla\mathcal{L}_{\text{vol}}(\theta_{t}))$
26:     if $\|d_{t}\|<\epsilon$ or $t=T_{\text{max}}$ then
27:         break
28:     end if
29:end for
30:return $\theta_{t}$ as $\theta^{*}$
IV EXPERIMENTS
Figure 4: The left, middle, and right subfigures present 2D slices of the learned inner safe sets obtained by the PCBF and ICCBF methods corresponding to the $x-y$, $x-v_{x}$, and $\psi-\omega_{z}$ planes. All other states are set to zero in each slice. In the left and middle subfigures, the obstacle’s projection in the corresponding 2D plane is indicated by black dashed lines, representing regions unsafe for the quadrotor.
In this section, we evaluate the performance of the PCBF algorithm. First, we use a 2D inverted pendulum example, where the low dimensionality allows HJ Reachability to compute the infinite-time viability kernel as the ground truth for the largest inner safe set. Comparing PCBF results with HJ validates its effectiveness. Next, we extend to a 12-dimensional quadrotor obstacle avoidance scenario and run simulations to assess PCBF’s performance in this more complex setting. Neural CBFs trained with linear combination loss (LCCBF) serve as our baseline. All neural networks are trained until convergence and share the same architecture (3 layers with 512, 256, 128 neurons and tanh activations).
IV-A Inverted Pendulum
The left subfigure of Figure 3 shows a strong overlap between the inner safe set generated by PCBF and the infinite-time viability kernel from HJ reachability. Almost all states in the PCBF inner safe set fall within the viability kernel, demonstrating that PCBF effectively maximizes the safe set while ensuring safety. In contrast, the LCCBF-generated set shows significant discrepancies, with many states outside the kernel, leading to potential safety violations in practice.
We visualized the training process of PCBF and LCCBF in the right subfigure in Figure 3. Initially, LCCBF effectively approximates the boundary of the largest inner safe set, showing that the linear combination loss quickly captures the safe region’s shape. However, as training progresses, LCCBF becomes unstable, with the learned inner safe set fluctuating unpredictably and failing to converge. In contrast, PCBF shows steady improvement as the volume of the learned safe set consistently growing
We also present the training times for different methods in Table I. In the sampling method column, "Boundary" refers to the sampling approach used in [6], while "Gaussian" represents the method proposed in this study. The results demonstrate a significant improvement in training efficiency with our new approach.
@@ IV-B Quadrotor
The state of the quadrotor involves its position $x,y,z$, velocity $v_{x},v_{y},v_{z}$, Euler angles $\phi,\theta,\psi$, and angular velocity $\omega_{x},\omega_{y},\omega_{z}$. The state dimension is 12 and the control dimension is 4. The system dynamics and input constraint can be found in [20]. A rectangular obstacle is placed in the scenario and the safe set is defined such that the quadrotor must maintain a minimum distance from the obstacle. Specifically, the safe set is $x\leq-1.5$ or $x\geq-0.75$ or $y\leq-0.375$ or $y\geq 0.375$.
@@
Figure 6: Minimum distance to the rectangle obstacle with and without the PCBF safety filter.
We conducted simulations using RotorPy [21] in a windless environment with a refresh rate of 300 Hz. The nominal policy is based on the control algorithm proposed by [22]. When the PCBF safety filter is activated, we solve the CBF-QP problem 5 to derive a safe control policy. Figure 5 illustrates the simulation results, highlighting the quadrotor trajectories with and without the PCBF safety filter. The trajectory without the safety filter collides with an obstacle, while the trajectory using the safety filter successfully avoids it.
Figure 6 presents the minimum distance between the quadrotor and the rectangular obstacle over time, comparing the scenarios with and without the safety filter. The red dashed line represents the minimum safe distance of 0.25 meters. When the safety filter is not applied, the quadrotor approaches the obstacle dangerously close, breaching the minimum safe distance. In contrast, with the safety filter enabled, the quadrotor maintains a safe distance throughout the simulation.
V CONCLUSIONS
In this paper, we introduce the Pareto Control Barrier Function (PCBF) framework to maximize the inner safe set under input constraints. Using a Pareto multi-task learning approach, we balanced safety and set size, with PCBF significantly outperforming LCCBF methods. Comparisons with HJ reachability for the inverted pendulum showed that PCBF closely approximates the largest inner safe set. Simulations with the quadrotor system further demonstrated PCBF’s effectiveness in complex, high-dimensional environments. Future work will focus on adapting the PCBF framework to handle uncertainties in system dynamics and external disturbances.
@@ V-A Proof of lemma 3
Let $\theta^{*}$ be the solution to problem 19. Suppose $\theta^{*}$ is not Pareto optimal. Then there exists a Pareto optimal $\hat{\theta}$ such that:
@@
with at least one of the inequalities being strict. However, since $\mathcal{L}_{\text{feas}}(\theta^{*})=0$, we must have $\mathcal{L}_{\text{feas}}(\hat{\theta})=0$ as well. This means $\hat{\theta}$ is also a feasible solution to problem 19. If $\mathcal{L}_{\text{vol}}(\hat{\theta})<\mathcal{L}_{\text{vol}}(\theta^{*})$, then $\hat{\theta}$ would be a better solution to problem 19 than $\theta^{*}$, which contradicts the assumption that $\theta^{*}$ is the solution to this problem. If $\mathcal{L}_{\text{vol}}(\hat{\theta})=\mathcal{L}_{\text{vol}}(\theta^{*})$, then $\hat{\theta}$ is also an optimal solution to problem 19, and thus satisfies the conditions of the lemma.
@@ V-B Analytical solution for PCBF subproblems
As problem (25) is a QP problem, we can apply the KKT conditions:

### v2

@@ III-C Pareto Control Barrier Function (PCBF)
Figure 3: The inner safe sets of the inverted pendulum problem obtained by the LCCBF (blue solid line) and PCBF (orange solid line) methods, alongside the infinite-time viability kernel (black dashed line) computed using HJ reachability. The region enclosed by the black dashed line represents the true safe states.
@@
Algorithm 1 Pareto Control Barrier Function (PCBF)
1:function SolveBaseSubproblem($\mathcal{L}_{1},\mathcal{L}_{2}$) $\triangleright$ See Appendix V-B $\lambda\leftarrow\max\left(\min\left(\frac{(\nabla\mathcal{L}_{2}-\nabla\mathcal{L}_{1})^{T}\nabla\mathcal{L}_{2}}{\|\nabla\mathcal{L}_{1}-\nabla\mathcal{L}_{2}\|_{2}^{2}},1\right),0\right)$
2:     return $-\lambda\nabla\mathcal{L}_{1}-(1-\lambda)\nabla\mathcal{L}_{2}$
3:end function
4:Initial parameter $\theta_{0}\in\mathbb{R}^{p}$, learning rate $\eta>0$, maximum iterations $T\in\mathbb{N}$, hyperparameters $\beta,\epsilon_{\text{ub}},\epsilon_{\text{lb}}$, regularization weight $\gamma>0$, safe state $x_{e}$, dataset $\mathbb{X}_{\mathcal{N}}$.
5:for $t=0,1,\ldots,T-1$ do
6:     $\mathcal{L}_{f}\leftarrow\mathcal{L}_{\text{feas}}(\theta_{t})$, $\mathcal{L}_{v}\leftarrow\mathcal{L}_{\text{vol}}(\theta_{t})$
7:     if $\mathcal{L}_{v}>\beta\mathcal{L}_{f}+\epsilon_{\text{ub}}$ then
8:         $d_{t}\leftarrow\textsc{SolveBaseSubproblem}(\mathcal{L}_{v}-\beta\mathcal{L}_{f},\mathcal{L}_{v})$triangleright Eq. (37), (38)
9:     else if $\beta\mathcal{L}_{f}+\epsilon_{\text{lb}}\leq\mathcal{L}_{v}\leq\beta\mathcal{L}_{f}+\epsilon_{\text{ub}}$ then
10:         $d_{t}\leftarrow\textsc{SolveBaseSubproblem}(\mathcal{L}_{f},\mathcal{L}_{v})$triangleright Eq. (39), (40)
11:     else
12:         $d_{t}\leftarrow\textsc{SolveBaseSubproblem}(\beta\mathcal{L}_{f}-\mathcal{L}_{v},\mathcal{L}_{f})$triangleright Eq. (41), (42)
13:     end if
14:     $g_{t}\leftarrow d_{t}-\gamma\nabla\mathcal{L}_{\text{vol}}(\theta_{t})$triangleright Apply regularization
15:     $\theta_{t+1}\leftarrow\theta_{t}+\eta g_{t}$
16:end for
17:return $\theta_{T}$
Figure 4: The left, middle, and right subfigures present 2D slices of the learned inner safe sets obtained by the PCBF and LCCBF methods corresponding to the $x-y$, $x-v_{x}$, and $\psi-\omega_{z}$ planes. All other states are set to zero in each slice. In the left and middle subfigures, the obstacle’s projection in the corresponding 2D plane is indicated by black dashed lines, representing regions unsafe for the quadrotor.
IV SIMULATIONS
In this section, the performance of the PCBF algorithm is evaluated. First, a 2D inverted pendulum example is employed, where the low dimensionality allows HJ Reachability to compute the infinite-time viability kernel as the ground truth for the largest inner safe set. Comparison of the PCBF results with HJ validates its effectiveness. Next, simulations were conducted in a 12-dimensional quadrotor obstacle avoidance scenario to assess PCBF’s performance in this more complex environment. Neural CBFs trained with linear combination loss (LCCBF) serve as our baseline. All neural networks are trained until convergence and share the same architecture (3 layers with 256 neurons and tanh activations).
IV-A Inverted Pendulum
Figure 3 shows a strong overlap between the inner safe set generated by PCBF and the infinite-time viability kernel from HJ reachability. Almost all states in the PCBF inner safe set fall within the viability kernel, demonstrating that PCBF effectively maximizes the safe set while ensuring safety. In contrast, the LCCBF-generated set shows significant discrepancies, with many states outside the kernel, leading to potential safety violations in practice.
The training times of different methods are reported in Table I. In the sampling method column, "Boundary" refers to the sampling approach used in [6], while "Gaussian" represents the method proposed in this study. The results demonstrate a significant improvement in training efficiency with our new approach.
@@ IV-B Quadrotor
The state of the quadrotor involves its position $x,y,z$, velocity $v_{x},v_{y},v_{z}$, Euler angles $\phi,\theta,\psi$, and angular velocity $\omega_{x},\omega_{y},\omega_{z}$. The state dimension is 12 and the control dimension is 4. The system dynamics and input constraints remain the same as in [21]. A rectangular obstacle is placed in the scenario and the safe set is defined such that the quadrotor must maintain a minimum distance from the obstacle. Specifically, the safe set is $x\leq-1.5$ or $x\geq-0.75$ or $y\leq-0.375$ or $y\geq 0.375$.
Simulations are conducted using RotorPy [22] in a windless environment with a refresh rate of 300 Hz. The nominal policy is based on the control algorithm proposed by [23]. When the PCBF safety filter is activated, problem 4 is solved to derive a safe control policy. Figure 5 illustrates the simulation results, highlighting the quadrotor trajectories with and without the PCBF safety filter. The trajectory without the safety filter collides with an obstacle, while the trajectory using the safety filter successfully avoids it.
V CONCLUSIONS
@@
In this paper, the Pareto Control Barrier Function (PCBF) algorithm is introduced to maximize the inner safe set under input constraints. The PCBF algorithm significantly outperforms traditional LCCBF methods. Comparisons with HJ reachability for the inverted pendulum showed that PCBF closely approximates the largest inner safe set. Simulations with the quadrotor system further demonstrated PCBF’s effectiveness in complex, high-dimensional environments. Future work will focus on adapting the PCBF framework to handle uncertainties in system dynamics and external disturbances.
@@ V-A Proof of lemma 2
Let $\theta^{*}$ be the solution to problem 15. Suppose $\theta^{*}$ is not Pareto optimal. Then there exists a Pareto optimal $\hat{\theta}$ such that:
@@
with at least one of the inequalities being strict. However, since $\mathcal{L}_{\text{feas}}(\theta^{*})=0$, we must have $\mathcal{L}_{\text{feas}}(\hat{\theta})=0$ as well. This means $\hat{\theta}$ is also a feasible solution to problem 15. If $\mathcal{L}_{\text{vol}}(\hat{\theta})<\mathcal{L}_{\text{vol}}(\theta^{*})$, then $\hat{\theta}$ would be a better solution to problem 15 than $\theta^{*}$, which contradicts the assumption that $\theta^{*}$ is the solution to this problem. If $\mathcal{L}_{\text{vol}}(\hat{\theta})=\mathcal{L}_{\text{vol}}(\theta^{*})$, then $\hat{\theta}$ is also an optimal solution to problem 15, and thus satisfies the conditions of the lemma.
@@ V-B Analytical solution for PCBF subproblems
As problem (21) is a QP problem, we can apply the Karush–Kuhn–Tucker (KKT) conditions:
@@
    $d_{t}=(\lambda+\lambda\beta-1)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\lambda\nabla\mathcal{L}_{\text{vol}}(\theta_{t})$ (37)
    where
    $(\displaystyle\lambda=\max\Bigg{(}\min\Bigg{(}$$\frac{((1+\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\nabla\mathcal{L}_{\text{vol}}(\theta_{t}))^{T}}{\|(1+\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\nabla\mathcal{L}_{\text{vol}}(\theta_{t})\|_{2}^{2}}$ (38)
    $\cdot\nabla\mathcal{L}_{\text{feas}}(\theta_{t}),\,1\Bigg{)},0\Bigg{)}$
  2. 2.
    $d_{t}=-\lambda\nabla\mathcal{L}_{\text{vol}}(\theta_{t})-(1-\lambda)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})$ (39)
    where
    $(\displaystyle\lambda=\max\Bigg{(}\min\Bigg{(}$$\frac{(\nabla\mathcal{L}_{\text{feas}}(\theta_{t})-\nabla\mathcal{L}_{\text{vol}}(\theta_{t}))^{T}\nabla\mathcal{L}_{\text{feas}}(\theta_{t})}{\|\nabla\mathcal{L}_{\text{vol}}(\theta_{t})-\nabla\mathcal{L}_{\text{feas}}(\theta_{t})\|_{2}^{2}},$ (40)
@@
  3. 3.
    $d_{t}=(\lambda-\lambda\beta-1)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})+\lambda\nabla\mathcal{L}_{\text{vol}}(\theta_{t})$ (41)
    where
    $(\displaystyle\lambda=\max\Bigg{(}\min\Bigg{(}$$\frac{(\nabla\mathcal{L}_{\text{vol}}(\theta_{t})+(1-\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t}))^{T}}{\|\nabla\mathcal{L}_{\text{vol}}(\theta_{t})+(1-\beta)\nabla\mathcal{L}_{\text{feas}}(\theta_{t})\|_{2}^{2}}$ (42)
    $\cdot\nabla\mathcal{L}_{\text{feas}}(\theta_{t}),\,1\Bigg{)},0\Bigg{)}$